find the least value of n for which the sum of the first n odd natural numbers exceeds 5000.


please give a very comprehensive answer easy to grasp.Thanks.

Answer by mananth(16946)

n odd numbers
1,3,5,7,9
Sum >5000
We will first find n for Sum = 5000
Sn = n/2(2a+(n-1)d)
a=1,d=2,Sn =5000
5000= n/2(2*1+(n-1)*2)
10000=n+(n-1)*2
10000=n+2n-2
10002=3n
n=3334
so for n =3334 Sum = 5000
SO for n=3335 Sum >5000